# Harmonic $1$-forms on real loci of Calabi-Yau manifolds

### v1

@@ 1 Introduction
[…]
@@
    There currently exists only one conjectural example for a Calabi-Yau manifold with the required $1$-form in the literature, namely [Joyce and Karigiannis, 2021, Section 7.5], based on the SYZ conjecture. Recent progress in proving the SYZ conjecture surveyed in [Li, 2022] does not support this conjectural example. Our numerical study suggests one new conjectural example.
@@
Acknowledgments. The second author wishes to thank Oliver Gäfvert, Andre Lukas, Lorenzo la Porta, Simon Salamon, and Mehdi Yazdi for helpful conversations. The third author would like to thank Simon Salamon, Challenger Mishra and Justin Tan for their hospitality during his visit at King’s College London and University of Cambridge, where part of this work was carried out. He also thanks James Halverson, Edward Hirst, Damián Mayorga Peña, and his supervisor Fabian Ruehle for useful discussion. The third author acknowledges support from the National Science Foundation under Cooperative Agreement PHY-2019786 (the NSF AI Institute for Artificial Intelligence and Fundamental Interactions).
@@ 2.2 $G_{2}$-orbifolds and their resolutions
The first examples of compact manifolds with holonomy $G_{2}$ were constructed in [Joyce, 1996]. This manifolds were obtained as the resolution of orbifolds $T^{7}/\Gamma$, where $\Gamma$ is a group of involutions acting on $T^{7}$.
@@
We denote $L:=\operatorname{fix}(\Gamma)$. By [Joyce, 2000, Prop. 10.8.1], the submanifold $L$ is a $3$-dimensional, three-dimensional, and an associative submanifold. The $G_{2}$-manifold $N$ is constructed by gluing an Eguchi-Hanson space into $M/\Gamma$ over every point in $L$. The construction requires a harmonic $1$-form $\lambda\in\Omega^{1}(L)$ which defines the size of the glued in Eguchi-Hanson space. Crucially, the construction only works if $\lambda$ is nowhere 00. It is an open problem to define a suitable $G_{2}$-structure around a zero of $\lambda$, see [Joyce and Karigiannis, 2021, Section 8, point (vi)].
@@ 3.1 Constructing nowhere vanishing harmonic $1$-forms from approximate harmonic $1$-forms
Let $(M,J,\omega,g)$ be a Calabi-Yau manifold with Calabi-Yau metric $g_{\text{CY}}$ and $L=\operatorname{fix}(\sigma)$ for an anti-holomorphic involution $\sigma:M\rightarrow M$. It is usually easy to compute the first Betti number of $L$, and if this is bigger than 00, then it is known to admit a harmonic $1$-form. However, it has been lamented above that it is difficult to check if such a $1$-form is nowhere vanishing.
@@
    Without explicitly computing $g_{\text{CY}}$, one can estimate the difference $\left|\left|g-g_{\text{CY}}\right|\right|_{C^{1,\alpha}}$ by computing a certain loss for $g$. This is explained in Corollary 3.5.
@@
    If the difference $\left|\left|g-g_{\text{CY}}\right|\right|_{C^{1,\alpha}}$ is sufficiently small, and there exists a harmonic $1$-form for $g$ that is sufficiently bounded from below, then that implies that also $g_{\text{CY}}$ must admit a harmonic $1$-form that is nowhere vanishing. This is a consequence of Proposition 3.6.
@@
if $(e^{f}-1)\in C^{1,\alpha}(M)$ has mean zero and satisfies $\left|\left|e^{f}-1\right|\right|_{C^{1,\alpha}}\leq\delta$, then the equation
@@
has a solution satisfying $\left|\left|\phi\right|\right|_{C^{3,\alpha/2}}\leq\epsilon$.
@@
The linearisation of Eq. 3.2 is $\phi\mapsto*\Delta\phi$. Assume the operator norm of the operator $\Delta^{-1}:C^{1,\alpha}_{0}(M)\rightarrow C^{3,\alpha}_{0}(M)$, mapping mean-zero functions to mean-zero functions, satisfies $\left|\left|\Delta^{-1}\right|\right|\leq C$. Define a sequence $\phi_{k}$ as follows: let $\phi_{0}=0$, and let $\phi_{k+1}$ be the unique mean-zero solution to
@@
(Note that this equation is essentially just Eq. 3.2 reordered so that the linear term is isolated on the left hand side.) For simplicity, write $K$ for a number satisfying $|\omega^{m}|\leq K$ pointwise everywhere. Let $\delta$ s.t. $4\delta^{2}C^{2}K^{2}m<\delta$. Then we can prove via induction that $\left|\left|\phi_{k}\right|\right|_{C^{3,\alpha}}\leq 2\delta CK$ for all $k\geq 0$, because:
$\left|\left|\phi_{k+1}\right|\right|_{C^{3,\alpha}}$
$\leq C\left|\left|\Delta\phi_{k+1}\right|\right|_{C^{1,\alpha}}$
@@
By the Arzelà-Ascoli theorem, we therefore have that $\phi_{k}$ has a subsequence converging in $C^{3,\alpha/2}$. By taking the limit in Eq. 3.3, we see that this limit solves Eq. 3.2. And by taking $\delta\leq\frac{\epsilon}{2CK}$ we have that $\left|\left|\phi_{k}\right|\right|_{C^{3,\alpha}}\leq 2\delta CK\leq\epsilon$, proving the claim. ∎
@@
Let $f\in C^{\infty}(M)$ be defined by $e^{f}\omega^{m}=\Omega\wedge\overline{\Omega}$. If $\left|\left|e^{f}-1\right|\right|_{C^{1,\alpha}}\leq\delta$, then $\left|\left|g-g_{\text{CY}}\right|\right|_{C^{1,\alpha/2}}\leq\epsilon$.
@@
$\left|\left|\lambda\right|\right|_{L^{2},g}=1,\text{ and }\min_{x\in L}|\lambda|_{g}(x)=\epsilon.$
There exists $\eta\in\Omega^{0}(L)$, unique up to constants, such that $\widetilde{\lambda}=\lambda+\operatorname{d}\!\eta$ is harmonic with respect to $h$. If $\left|\left|g-h\right|\right|_{C^{1},g}\leq\delta$, then $\widetilde{\lambda}$ has no zeros.
@@
Assume $\min_{x\in L}|\lambda|_{g}(x)=\epsilon$ and $\left|\left|g-h\right|\right|_{C^{1},g}\leq\delta$ with $\delta$ to be determined later. The operators $\Delta$ and $\operatorname{d}\!^{*}$ depend on the metric, and we denote the operators corresponding to $g$ and $h$ by $\Delta_{g}$, $\operatorname{d}\!^{*}_{g}$ and $\Delta_{h}$, $\operatorname{d}\!^{*}_{h}$ respectively.
@@
$\left|\left|\operatorname{d}\!^{*}_{h}\lambda\right|\right|_{L^{p},h}=\left|\left|(\operatorname{d}\!^{*}_{h}-d^{*}_{g})\lambda\right|\right|_{L^{p},h}\leq\left|\left|h-g\right|\right|_{C^{1},h}\cdot\left|\left|\lambda\right|\right|_{L^{p},h}\lesssim\delta.$ (3.7)
Here, in the last step we used that $\left|\left|g-h\right|\right|_{C^{1},g}\leq\delta$ implies $\left|\left|g-h\right|\right|_{C^{1},h}\lesssim\delta$ if $\delta<\frac{1}{2}$, which we assumed without loss of generality. We also used $\left|\left|\lambda\right|\right|_{L^{p},h}\lesssim\left|\left|\lambda\right|\right|_{L^{2},h}\lesssim\left|\left|\lambda\right|\right|_{L^{2},g}=1.$ In this, the first inequality is true because $\operatorname{Ker}(\Delta_{g})$ is a finite-dimensional vector space, so on it all norms are equivalent; the second inequality follows from $\left|\left|g-h\right|\right|_{C^{1},h}\lesssim\delta$ with the assumption that $\delta<\frac{1}{2}$. Now let $\eta\in\Omega^{0}(L)$ be the unique function with mean zero satisfying the equation
@@
$\left|\left|\operatorname{d}\!\eta\right|\right|_{C^{0},h}\leq\left|\left|\eta\right|\right|_{C^{1,\alpha},h}\lesssim\left|\left|\eta\right|\right|_{L^{p}_{2},h}\lesssim\left|\left|\Delta_{h}\eta\right|\right|_{L^{p},h}=\left|\left|\operatorname{d}\!^{*}_{h}\lambda\right|\right|_{L^{p},h}\lesssim\delta.$
@@
$\left|\left|\operatorname{d}\!\eta\right|\right|_{C^{0},g}\lesssim\left|\left|\operatorname{d}\!\eta\right|\right|_{C^{0},h}\lesssim\delta,$
where we again assumed $\delta<\frac{1}{2}$. Spelled out, this means that there exists a constant $C$ depending only on $h$ and the assumption that $\delta<\frac{1}{2}$ but not on $g$ or $\delta$, such that $\left|\left|\operatorname{d}\!\eta\right|\right|_{C^{0},g}\leq C\delta$. Now, if $\delta<\epsilon/C$, then
@@
Let $f\in C^{\infty}(M)$ be defined by $e^{f}\omega^{m}=\Omega\wedge\overline{\Omega}$. If $\left|\left|e^{f}-1\right|\right|_{C^{1,\alpha}}\leq\delta$ and $\lambda\in\Omega^{1}(L)$ is harmonic with respect to $g$ and
$\left|\left|\lambda\right|\right|_{L^{2},g|_{L}}=1,\text{ and }\min_{x\in L}|\lambda|_{g|_{L}}(x)=\epsilon$
@@ 3.2.2 The quintic Quintic with $b_{1}(L)=1$
    Given $D=Z(f)$ we obtain the singular quintic $Q^{*}=Z((x_{0}^{2}+\dots+x_{4}^{2})f)$ which is bijective to $D$. A suitable smoothing $Q$ of $Q^{*}$ will then satisfy $b_{1}(Q)=1$.
@@
It was shown in [Krasnov, 2006] that the space of smooth cubic surfaces in $\mathbb{RP}^{4}$ consists of $8$ chambers, separated by $14$ walls. The exact arrangement of chambers and walls is given in [Krasnov, 2009, p.843]. All walls are sets of singular cubics with a single ordinary double point, and the walls can be distinguished using information coming from the ordinary double point.
We now define a singular cubic $D^{*}$ in one of these walls, namely $\mathscr{D}^{3}_{1}(1)$ in the notation of [Krasnov, 2009]. The notation $\mathscr{D}^{3}_{1}(1)$ means that the bilinear form defining the ordinary double point of $D^{*}$ has signature $(3,1)$, and that the tautological curve consisting of all lines in $\mathbb{P}(T_{a}\mathbb{RP}^{4})$ that lie completely in $D^{*}$ has exactly one connected component.
@@
Note that in this case it is known that every harmonic $1$-form on $D$ must have at least two zeros because of the following proposition:
@@
Every closed $1$-form on $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ must have at least two zeros, and it must have an even number of zeros.
@@
The manifold $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ is irreducible in the sense of [Jaco, 1980, p.18]. By [Jaco, 1980, VI.7. Lemma], $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ is not Seifert fibered, in particular it is not a fibration over $S^{1}$. By [Tischler, 1970], $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ does not admit a nowhere vanishing closed $1$-form. In particular, a harmonic $1$-form must have at least one zero. As $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ is orientable, its Euler characteristic $\chi(\mathbb{R}P^{3}\#(S^{1}\times S^{2}))$ is zero. By the Poincaré–Hopf theorem (see e.g. [Milnor, 1965, p.35]), this implies that a harmonic $1$-form must have an even number of zeros. ∎
@@
Define the quintic polynomial $g^{*}:=(x_{0}^{2}+\dots+x_{4}^{2})f$. Notice first that the factor $(x_{0}^{2}+\dots+x_{4}^{2})$ is nowhere vanishing on $\mathbb{RP}^{4}$. Thus, $Z_{\mathbb{R}}(f)=Z_{\mathbb{R}}(g^{*})\subset\mathbb{RP}^{4}$. The real locus of the complex variety $X^{*}:=Z(g^{*})$ is smooth, but for our application we require an algebraic variety that is smooth as a complex variety. The variety $X^{*}\subset\mathbb{CP}^{4}$ is reducible, so in particular singular, but a small generic perturbation of it is smooth. As $Z_{\mathbb{R}}(g^{*})\subset\mathbb{RP}^{4}$ is smooth, any small perturbation of it does not change its diffeomorphism type. Altogether, we have shown the following:
@@
where $\epsilon<\epsilon_{0}$ is a non-zero real number near 00 and $P\in\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$ is a generic homogeneous polynomial of degree $5$, defines a smooth algebraic variety in $\mathbb{CP}^{4}$, and its real locus has first Betti number $b^{1}=1$.
@@
Fix homogeneous degree $5$ polynomials $g^{*},P$ such that $Z_{\mathbb{R}}(g^{*})\subset\mathbb{RP}^{4}$ is smooth. Let $U\subset S^{4}$ be a neighbourhood of $Z_{\mathbb{R}}(g^{*})\cap S^{4}\subset\mathbb{R}^{5}$. Let $\widetilde{g}(x):=g^{*}|_{S^{4}},\widetilde{P}(x):=P|_{S^{4}}\in C^{\infty}(S^{4})$. Define
$m:=\min_{x\in U}|\nabla\widetilde{g}|,\quad k:=\min_{x\in S^{4}\setminus U}|\widetilde{g}|.$
If $\left|\left|\widetilde{P}\right|\right|_{C^{0}}<k$ and $\left|\left|\nabla\widetilde{P}\right|\right|_{C^{0}}<m$, then $X:=Z_{\mathbb{R}}(g^{*}+\epsilon P)\subset\mathbb{RP}^{4}$ is smooth for all $\epsilon\in[0,1)$.
@@
The condition $\left|\left|\widetilde{P}\right|\right|_{C^{0}}<k$ ensures that $\widetilde{g}+\epsilon\widetilde{P}$ has no zeros outside of $U$, i.e. in particular the variety $Z(\widetilde{g}+\epsilon\widetilde{P})$ has no singular points outside of $U$. The condition $\left|\left|\nabla\widetilde{P}\right|\right|_{C^{0}}<m$ ensures that $\nabla(\widetilde{g}+\epsilon\widetilde{P})\neq 0$ on $U$, in particular the variety $Z(\widetilde{g}+\epsilon\widetilde{P})$ has no singular points in $U$. ∎
@@
Using numerical minimisation together with Proposition 3.12 suggests that also $Z(g^{*}+\epsilon P)$ is smooth for every $\epsilon\leq\frac{27}{1000}$, however we remark that due to the use of numerical minimisation we have no rigorous proof of that fact. Because of this, we cannot guarantee that the real locus is diffeomorphic to $Z(f)\subset\mathbb{RP}^{4}$. Using the persistent homology tool from [Gäfvert, 2022, Di Rocco et al., 2022] suggests that $b_{1}(Z(g^{*}+\epsilon P)=1$, though also here we did no provably correct calculation of this Betti number (which would be available using the same tool with extra effort).

### v2

@@ 1 Introduction
For some applications it would be desirable to have information about the metric. In Physics, this is needed to compute particle masses and couplings, and very often approximate Calabi-Yau metrics are used to approximately compute these quantities, see e.g. [Ashmore and Ruehle, 2021, Constantin et al., 2024, Butbaia et al., 2024]. In this article, we focus on the following application in differential geometry: in [Joyce and Karigiannis, 2021] a new construction method for $G_{2}$-manifolds, which are Ricci-flat Riemannian manifolds in dimension $7$, was explained. Contrary to Calabi-Yau manifolds, not many constructions for these manifolds are known. However, they are important objects in M-theory and differential geometry, and new examples would help to generate conjectures tackling open questions in the field, for example which smooth $7$-manifolds can be $G_{2}$-manifolds. The construction in [Joyce and Karigiannis, 2021] requires as input a Calabi-Yau manifold $M$ of complex dimension $3$ whose real locus $L\subset M$ (of real dimension $3$) admits a harmonic nowhere vanishing $1$-form $\lambda\in\Omega^{1}(L)$. Here, harmonic means harmonic with respect to the elusive Calabi-Yau metric. So far, the nowhere vanishing condition could not be checked in any non-trivial example. (It could be checked on the complex $3$-dimensional manifold $T^{2}\times\text{K3}$, where $T^{2}$ denotes the torus of real dimension $2$.) The construction method gives quite an explicit description of the Ricci-flat metric on the $7$-manifold. This would allow approximately computing some quantities in M-theory on these examples, paralleling the previously described computations on Calabi-Yau manifolds from the physics literature.
@@
    There currently exists only one conjectural example for a Calabi-Yau manifold with the required $1$-form in the literature, namely [Joyce and Karigiannis, 2021, Section 7.5]. Our numerical study suggests one new conjectural example.
@@
Acknowledgments. The second author wishes to thank Oliver Gäfvert, Michel van Garrel, Andre Lukas, Lorenzo la Porta, Simon Salamon, Calum Spicer, and Mehdi Yazdi for helpful conversations. The third author would like to thank Simon Salamon, Challenger Mishra and Justin Tan for their hospitality during his visit at King’s College London and University of Cambridge, where part of this work was carried out. He also thanks James Halverson, Edward Hirst, Damián Mayorga Peña, and his supervisor Fabian Ruehle for useful discussion. The third author acknowledges support from the National Science Foundation under Cooperative Agreement PHY-2019786 (the NSF AI Institute for Artificial Intelligence and Fundamental Interactions). The fourth author would like to thank Simon Donaldson for helpful conversations, and to acknowledge support from the Simons Center for Geometry and Physics. The authors thank the anonymous referee for their helpful comments.
@@ 2.2 $G_{2}$-orbifolds and their resolutions
The first examples of compact manifolds with holonomy $G_{2}$ were constructed in [Joyce, 1996]. These manifolds were obtained as the resolution of orbifolds $T^{7}/\Gamma$, where $\Gamma$ is a group of involutions acting on $T^{7}$.
@@
We denote $L:=\operatorname{fix}(\Gamma)$. By [Joyce, 2000, Prop. 10.8.1], the submanifold $L$ is a three-dimensional associative submanifold. The $G_{2}$-manifold $N$ is constructed by gluing an Eguchi-Hanson space into $M/\Gamma$ over every point in $L$. The construction requires a harmonic $1$-form $\lambda\in\Omega^{1}(L)$ which defines the size of the glued in Eguchi-Hanson space. Crucially, the construction only works if $\lambda$ is nowhere 0. It is an open problem to define a suitable $G_{2}$-structure around a zero of $\lambda$, see [Joyce and Karigiannis, 2021, Section 8, point (vi)].
@@ 3.1 Constructing nowhere vanishing harmonic $1$-forms from approximate harmonic $1$-forms
Let $(M,J,\omega,g)$ be a Calabi-Yau manifold with Calabi-Yau metric $g_{\text{CY}}$ and $L=\operatorname{fix}(\sigma)$ for an involution $\sigma:M\rightarrow M$ that is anti-holomorphic. It is usually easy to compute the first Betti number of $L$, and if this is bigger than 0, then it is known to admit a harmonic $1$-form. However, it has been lamented above that it is difficult to check if such a $1$-form is nowhere vanishing.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.2.2 The quintic Quintic with $b_{1}(L)=1$
    Given $D=Z_{\mathbb{R}}(f)$ we obtain the singular quintic $Q^{*}=Z_{\mathbb{R}}((x_{0}^{2}+\dots+x_{4}^{2})f)$ which is bijective to $D$. A suitable smoothing $Q$ of $Q^{*}$ will then still satisfy $b_{1}(Q)=1$.
@@
It was shown in the work [Krasnov, 2006] that the space of smooth cubic surfaces in $\mathbb{RP}^{4}$ consists of $8$ chambers, separated by $14$ walls. The exact arrangement of chambers and walls is given in [Krasnov, 2009, p.843]. All walls are sets of singular cubics with a single ordinary double point, and the walls can be distinguished using information coming from the ordinary double point.
We now define a singular cubic $D^{*}$ in one of these walls, namely $\mathscr{D}^{3}_{1}(1)$ in the notation from the article [Krasnov, 2009]. The notation $\mathscr{D}^{3}_{1}(1)$ means that the bilinear form defining the ordinary double point of $D^{*}$ has signature $(3,1)$, and that the tautological curve consisting of all lines in $\mathbb{P}(T_{a}\mathbb{RP}^{4})$ that lie completely in $D^{*}$ has exactly one connected component.
@@
Note that in this case it is known that every harmonic $1$-form on $D$ must have at least one zero because of the following proposition:
@@
Every closed $1$-form on $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ must have at least one zero. If its zeros are non-degenerate (i.e. near each zero it is locally of the form $\operatorname{d}\!f$ for a function $f$ with non-degenerate Hessian), then it must have an even number of zeros.
@@
The manifold $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ is reducible in the sense of [Jaco, 1980, p.18]. By [Jaco, 1980, VI.7. Lemma], $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ is not Seifert fibered, in particular it is not a fibration over $S^{1}$. By [Tischler, 1970], $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ does not admit a nowhere vanishing closed $1$-form. In particular, a harmonic $1$-form must have at least one zero. As $\mathbb{R}P^{3}\#(S^{1}\times S^{2})$ is orientable, its Euler characteristic $\chi(\mathbb{R}P^{3}\#(S^{1}\times S^{2}))$ is zero. By the Poincaré–Hopf theorem (see e.g. [Milnor, 1965, p.35]), this implies that a harmonic $1$-form with non-degenerate zeros must have an even number of zeros. ∎
@@
Define the quintic polynomial $g^{*}:=(x_{0}^{2}+\dots+x_{4}^{2})f$. Notice first that the factor $(x_{0}^{2}+\dots+x_{4}^{2})$ is nowhere vanishing on $\mathbb{RP}^{4}$. Thus, $Z_{\mathbb{R}}(f)=Z_{\mathbb{R}}(g^{*})\subset\mathbb{RP}^{4}$. The real locus of the complex variety $X^{*}:=Z_{\mathbb{C}}(g^{*})$ is smooth, but for our application we require an algebraic variety that is smooth as a complex variety. The variety $X^{*}\subset\mathbb{CP}^{4}$ is reducible, so in particular singular, but a small generic perturbation of it is smooth. As $Z_{\mathbb{R}}(g^{*})\subset\mathbb{RP}^{4}$ is smooth, any small perturbation of it does not change its diffeomorphism type. Altogether, we have shown the following:
@@
where $\epsilon\in\mathbb{R}$ such that $0<\epsilon<\epsilon_{0}$ and $P\in\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$ is a generic homogeneous polynomial of degree $5$, defines a smooth algebraic variety in $\mathbb{CP}^{4}$, and its real locus has first Betti number $b^{1}=1$.
@@
Fix homogeneous degree $5$ polynomials $g^{*},P$ such that $Z_{\mathbb{R}}(g^{*})\subset\mathbb{RP}^{4}$ is smooth. Let $U\subset S^{4}$ be a neighbourhood of $Z_{\mathbb{R}}(g^{*})\cap S^{4}\subset\mathbb{R}^{5}$. Let $\widetilde{g}:=g^{*}|_{S^{4}},\widetilde{P}:=P|_{S^{4}}\in C^{\infty}(S^{4})$. Define
[…]
@@
[…]
@@
Using numerical minimisation together with Proposition 3.12 suggests that also $Z(g^{*}+\epsilon P)$ is smooth for every $\epsilon\leq\frac{27}{1000}$, however we remark that due to the use of numerical minimisation we have no rigorous proof of that fact. Because of this, we cannot guarantee that the real locus is diffeomorphic to $Z(f)\subset\mathbb{RP}^{4}$. Using the persistent homology tool from [Gäfvert, 2022, Di Rocco et al., 2022] suggests that $b_{1}(Z(g^{*}+\epsilon P)=1$, though also here we did no provably correct calculation of this Betti number (which would be available using the same tool but is computationally very expensive).